I dont understand how to work this problem..
9/n = 5 over 11/15

Answer by checkley77(12844)

9/n = 5 over 11/15
9/N=5/(11/15)
9/N=5*15/11
9/N=75/11
75N=99
N=99/75
N=33/25 ANS.